If sin A=6÷13 and sec B=13/6 then A+B is equal to
Show solution without calculator
150°
90°
60°
45°

Answer by greenestamps(13200)



I will assume these angles are less than 90 degrees each; otherwise there are an infinite number of values for the sum of the measures of A and B.

sinA = 6/13 and secB = 13/6 --> sinA = 6/13 and cosB = 6/13.

In the first quadrant, if sinA = cosB, then A and B are complementary angles, so A+B = 90 degrees.